OpenStudy (anonymous):

What is the line of symmetry for the parabola whose equation is y = -x^2 + x + 3? 1.) x=-1 2.) x= -1/2 3.) x= 1/2

OpenStudy (anonymous):

same as before, compute \[\huge -\frac{\color{blue}b}{2\color{red}a}\] with \[\large \color{red}a=\color{red}{-1},\color{blue}b=\color{blue}1\]

OpenStudy (anonymous):

So -1 divided by 2(-1) ? And that would be 3.) x= 1/2 ?

OpenStudy (anonymous):

yes it would!

OpenStudy (anonymous):

Thank you:)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

I just don't understand how you set it up

OpenStudy (anonymous):

i am not sure what else to say other than the axis of symmetry is \[\large x= -\frac{\color{blue}b}{2\color{red}a}\]for \[\large y= \color{red}ax^2+\color{blue}bx+c\]

OpenStudy (anonymous):

Thanks, again:)

OpenStudy (anonymous):

yw